OpenStudy (anonymous):

Which of the following is equivalent to3 2/5 divided by 4? 1 3/17 3 1/10 15/20 17/20

OpenStudy (anonymous):

\[\large 3 \frac{2}{5} \implies \frac{5 \times 3 + 2}{5} \implies \frac{17}{5}\]

mathslover (mathslover):

now divide that by 4

mathslover (mathslover):

that is : \[\large{\frac{17}{5*4}=\frac{17}{20}}\]

OpenStudy (anonymous):

\[\large \frac{\frac{17}{5}}{4} \implies \frac{17}{5 \times 4} = ??\]

OpenStudy (anonymous):

17/20

mathslover (mathslover):

good that is ur answer

OpenStudy (anonymous):

Well Done..

mathslover (mathslover):

d) 17/20 <----- ur answer @merp -_- :)

OpenStudy (anonymous):

oh i get thequestion now thanks guys :)